### v1

@@ 2.3 Difference-in-Differences
As is the case of the representation in Proposition 2.4, the two-way fixed effects estimand in Proposition 2.5 satisfies the representation in (1.1), with $X=G$, $W_{0}=D$, $w_{0}(X)=\mathbb{P}(D=1\mid G)$, $\tau_{0}(X)=\mathbb{E}[Y(1)-Y(0)\mid D=1,G]$, and the weight function $a(X)=a_{B}(G)\geq 0$. This weight function is derived in the proof of the proposition, and we show it is equivalent to equation (16) in Goodman-Bacon (2021) in Appendix D.
3 Causal Representation of Weighted Estimands
We now consider a class of weighted estimands that subsumes all our previous examples. In this section, we show necessary and sufficient conditions for an estimand of this class to have a causal representation as an average treatment effect over a population. Then, we compute the largest size of a population that is compatible with the weighted estimand. The size of this population relative to $\mathbb{P}(W_{0}=1)$ will be a measure of the internal validity of our estimand relative to the target estimand $\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$. The absolute size of this population will be a measure of the representativeness of the estimand with respect to the population distribution the data is drawn from.
Recall the earlier setting where we let $D\in\{0,1\}$ denote a binary treatment variable, and let $(Y(1),Y(0))$ denote the corresponding potential outcomes under treatment and control, respectively. Let $X\in\text{supp}(X)\subseteq\mathbb{R}^{d_{X}}$ denote a $d_{X}$-vector of covariates. We suppose that $(Y(1),Y(0),D,X)$ are drawn from a common population distribution $F_{Y(1),Y(0),D,X}$.
Let $W_{0}\in\{0,1\}$ be an indicator variable used to denote a subpopulation $\{W_{0}=1\}$. Throughout this paper, we assume that $\mathbb{P}(W_{0}=1)>0$, so that this subpopulation has a positive mass, which avoids technical issues associated with conditioning on zero-probability events. For example, this subpopulation can be the entire population by setting $W_{0}=1$ almost surely, the treated units by setting $W_{0}=D$, or units complying with binary instrument $Z$ by setting $W_{0}=\mathbbm{1}(D(1)>D(0))$. Let $\tau_{0}(X)=\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1,X]$ denote the conditional average treatment effect given $X$ in subpopulation $W_{0}$. This is defined for all values of $X$ such that $w_{0}(X)>0$.
Also recall the weighted estimands of equation (1.1):
$\mu(a,\tau_{0})$ $\equiv\frac{\mathbb{E}[a(X)w_{0}(X)\tau_{0}(X)]}{\mathbb{E}[a(X)w_{0}(X)]}=\frac{\mathbb{E}[a(X)\tau_{0}(X)\mid W_{0}=1]}{\mathbb{E}[a(X)\mid W_{0}=1]},$
where $w_{0}(X)=\mathbb{P}(W_{0}=1\mid X)$. [^3] While $\tau_{0}(X)$ is only defined for $w_{0}(X)>0$, we set $\tau_{0}(X)w_{0}(X)=0$ when $w_{0}(X)=0$.
The estimands we consider have the above representation and satisfy the following regularity conditions.
Assumption 3.1 (Regularity).
The first two restrictions are weak regularity assumptions that ensure the finiteness of the numerator of $\mu(a,\tau_{0})$. We rule out $\mathbb{E}[a(X)\mid W_{0}=1]=0$ since it implies the estimand does not exist. The estimand in (1.1) is unchanged if the sign of $a(X)$ is reversed, so $\mathbb{E}[a(X)\mid W_{0}=1]>0$ is a sign normalization.
3.1 Alternative Representations of Weighted Estimands
The weighted estimands of equation (1.1) can also be written as a weighted sum when $X$ is discrete, or an integral when $X$ is continuous. In the discrete case, let $\text{supp}(X)=\{x_{1},\ldots,x_{K}\}$, $W_{0}=1$ almost surely, and let $p_{k}\equiv\mathbb{P}(X=x_{k})>0$ for $k=1,\ldots,K$. Then,
$\mu(a,\tau_{0})$ $=\sum_{k=1}^{K}\omega_{k}\tau_{0}(x_{k})\qquad\text{ where }\qquad\omega_{k}=\frac{a(x_{k})p_{k}}{\sum_{l=1}^{K}a(x_{l})p_{l}},$ (3.1)
which are weights that sum to one. The representations in (1.1) and (3.1) are equivalent as we can obtain $a(x_{k})$ (up to scale) as the ratio $\omega_{k}/p_{k}$, and $\omega_{k}$ is defined as a function of $(a(\cdot),p_{\cdot})$ in equation (3.1) above.
From (3.1), we can see that $a(x_{k})$ being constant ensures $\omega_{k}=p_{k}$, or that the estimand is the ATE. Moreover,
$\frac{a(x_{k})}{a(x_{k^{\prime}})}=\frac{\omega_{k}}{\omega_{k^{\prime}}}\big{/}\frac{p_{k}}{p_{k^{\prime}}},$
which is the ratio of the relative weights of covariate cells $\{X=x_{k}\}$ and $\{X=x_{k^{\prime}}\}$ in the estimand $(\omega_{k}/\omega_{k^{\prime}})$ and in the population $(p_{k}/p_{k^{\prime}})$. The inequality $a(x_{k})>a(x_{k^{\prime}})$ indicates that covariate cell $\{X=x_{k}\}$ is overweighted relative to $\{X=x_{k^{\prime}}\}$, when compared to their relative weights in the population.
Alternatively, consider the case where $X$ is continuously distributed with density [^4] With respect to the Lebesgue measure. $f_{X}$. Still assuming $W_{0}=1$ almost surely, we can write

### v3

@@ 5.3   Difference-in-Differences
As is the case of the representation in Proposition 5.3, the two-way fixed effects estimand in Proposition 5.4 satisfies the representation in (1.1), with $X=G$, $W_{0}=D$, $\tau_{0}(X)=\mathbb{E}[Y(1)-Y(0)\mid D=1,G]$, and the weight function $a_{\text{TWFE,H}}(G)\geq 0$. This weight function is derived in Appendix E, with additional comparisons with the weights in Goodman-Bacon (2021) in Appendix H.
Causal Representation and Internal Validity of TWFE
We now consider the weights obtained in Proposition 5.4 under its assumptions. These weights are nonnegative and therefore Theorem 3.1 guarantees the existence of a causal representation for $\beta_{\text{TWFE}}$ uniformly in $\tau_{0}\in\mathcal{T}_{\text{all}}$. Using Theorem 4.1, the internal validity of $\beta_{\text{TWFE}}$ relative to target parameter $\mathbb{E}[Y(1)-Y(0)\mid D=1]$ is given by
$\overline{P}(a_{\text{TWFE,H}},D;\mathcal{T}_{\text{all}})=\frac{\mathbb{E}[a_{\text{TWFE,H}}(G)\mid D=1]}{\sup_{g\in\text{supp}(G\mid D=1)}a_{\text{TWFE,H}}(g)}$
Due to the absorbing nature of the treatment in our setting, all expressions involving the distribution of $D$ given $P$ or $G$ can be derived as a function of the marginal distribution of $G$. Therefore, $\overline{P}(a_{\text{TWFE,H}},D;\mathcal{T}_{\text{all}})$ depends only on $\{\mathbb{P}(G=g)\}_{g\in\{2,\ldots,T\}}$.
To give some intuition, consider the case where $T=3$ and therefore $G\in\{2,3,+\infty\}$. In this case, calculations yield that
$\overline{P}(a_{\text{TWFE,H}},D;\mathcal{T}_{\text{all}})$ $=\min\left\{\frac{2\mathbb{P}(G=2)+\omega\mathbb{P}(G=3)}{2\mathbb{P}(G=2)+\mathbb{P}(G=3)},\frac{\omega^{-1}2\mathbb{P}(G=2)+\mathbb{P}(G=3)}{2\mathbb{P}(G=2)+\mathbb{P}(G=3)}\right\},$
where $\omega=\frac{4-2\mathbb{P}(G=2)-4\mathbb{P}(G=3)}{2-2\mathbb{P}(G=2)-\mathbb{P}(G=3)}=\frac{a_{\text{TWFE,H}}(2)}{a_{\text{TWFE,H}}(3)}$. Therefore, $\beta_{\text{TWFE}}$ is perfectly representative of the ATT if and only if $\omega=1$, which occurs if and only if $\mathbb{P}(G=3)=2/3$. Thus, $\overline{P}(a_{\text{TWFE,H}},D;\mathcal{T}_{\text{all}})=1$ when $\mathbb{P}(G=3)=2/3$ and the internal validity of $\beta_{\text{TWFE}}$ declines as $|\mathbb{P}(G=3)-2/3|$ increases. This is due to the weight function $a_{\text{TWFE,H}}(g)$ being constant in $g$ if and only if $\mathbb{P}(G=3)=2/3$. As before, constant weights imply that the weighted estimand equals the average treatment effect over $\{W_{0}=1\}$.
6   Estimation and Inference
We now consider estimation and inference for our measures of internal validity and representativeness. We focus our attention on the case when $\mathcal{T}=\mathcal{T}_{\text{all}}$ and briefly discuss the case where $\mathcal{T}=\{\tau_{0}\}$ in Appendix C.
To measure internal validity, we seek to estimate
$\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})=\frac{\mathbb{E}[a(X)\mid W_{0}=1]}{a_{\max}}=\frac{\mathbb{E}[a(X)w_{0}(X)]}{\mathbb{E}[w_{0}(X)]\cdot a_{\max}}.$
Suppose we observe a random sample of size $n$, $\{(W_{i},X_{i})\}_{i=1}^{n}$, where $W_{i}$ is a set of variables needed to estimate $a(\cdot)$ and $w_{0}(\cdot)$. For example, under unconfoundedness we can let $W_{i}=D_{i}$ since the distribution of $(D,X)$ is sufficient to identify $a(\cdot)$; the outcome’s distribution does not affect $\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})$. In our instrumental variables examples, we let $W_{i}=(D_{i},Z_{i})$.
Assuming the existence of estimators for $a(\cdot)$ and $w_{0}(\cdot)$, we consider the following analog estimator of $\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})$:
$\widehat{\overline{P}}$ $\coloneqq\frac{\frac{1}{n}\sum_{i=1}^{n}\widehat{a}(X_{i})\widehat{w}_{0}(X_{i})}{\frac{1}{n}\sum_{i=1}^{n}\widehat{w}_{0}(X_{i})\cdot\max_{i:\widehat{w}_{0}(X_{i})>c_{n}}\widehat{a}(X_{i})}.$
Here $c_{n}$ is a tuning parameter that converges to 0 as $n$ diverges. We start by noting that we can estimate $\mathbb{E}[a(X)\mid W_{0}=1]$ via $\frac{\frac{1}{n}\sum_{i=1}^{n}\widehat{a}(X_{i})\widehat{w}_{0}(X_{i})}{\frac{1}{n}\sum_{i=1}^{n}\widehat{w}_{0}(X_{i})}$, which will be consistent under standard conditions on $\widehat{a}(\cdot)$ and $\widehat{w}_{0}(\cdot)$. Estimating $a_{\max}=\sup(\text{supp}(a(X)\mid W_{0}=1))$ is more delicate. In some of our examples, this supremum is known or can be bounded above without using data. For example, the OLS estimand under unconfoundedness has weights $a(X)=\operatorname{var}(D\mid X)$ which are bounded above by $1/4$. If $X$ is continuously distributed, then 1/2 may lie in the support of $p(X)$, and thus we may avoid the estimation of $a_{\max}$. The IV estimand of Section 5 has weights $a_{\text{IV}}(X)=\operatorname{var}(Z\mid X)$ which are similarly bounded above by $1/4$. Similarly, $a_{\text{2SLS}}(X)=|\operatorname{cov}(D,Z\mid X)|\leq 1/4$ by the Cauchy–Schwarz inequality. If knowledge of $a_{\max}$ is not assumed, but $\text{supp}(X\mid W_{0}=1)$ is known and $a(x)$ is continuous, [^4] Note that $a(x)$ is trivially continuous on finite support. then $\sup_{x\in\text{supp}(X\mid W_{0}=1)}\widehat{a}(x)$ will be consistent for $a_{\max}$ when $\widehat{a}(x)$ is consistent for $a(x)$ uniformly in $x\in\text{supp}(X\mid W_{0}=1)$. Many parametric and nonparametric estimators for $a(\cdot)$ satisfy this requirement.
In Appendix C, we prove the consistency of $\widehat{\overline{P}}$ and derive its limiting distribution. We also provide a step-by-step bootstrap algorithm that can be employed to conduct inference and prove its validity. This bootstrap approach is based on Fang and Santos (2019) and is nonstandard, but yields valid inferences even when $a_{\max}$ is estimated, as opposed to standard bootstrap approaches, such as the empirical bootstrap.
7   Empirical Application
In this section, we implement the proposed tools in an application to the effects of unilateral divorce laws in the U.S. on female suicide, as in Stevenson and Wolfers (2006). Between 1969 and 1985, 37 states (including D.C.) reformed their law by enabling each spouse to seek divorce without the other spouse’s consent. Stevenson and Wolfers (2006) argue that these “unilateral” or “no-fault” divorce laws reduced female suicide, domestic violence, and spousal homicide. The results on female suicide are also replicated by Goodman-Bacon (2021), whose analysis we follow here.
Our sample consists of 41 states observed over the 1964–1996 period. The outcome of interest is the state- and year-specific female suicide rate, as computed by the National Center for Health Statistics. The treatment is whether the state allowed unilateral divorce in a given year. Following Goodman-Bacon (2021), our sample omits Alaska and Hawaii. We also omit eight further states which had unilateral divorce laws preceding 1964 and are therefore always treated within our timeframe.
The conclusion changes, however, when we explicitly target the average treatment effect on the treated (ATT), that is, the average effect for the largest subpopulation for which such an effect is identified under standard assumptions. Using the approach of Wooldridge (2021), we obtain an estimate of –5.530 with a $p$-value of 0.138. The approach of Callaway and Sant’Anna (2021) produces an estimate of –10.220 and a $p$-value of 0.001. These estimates are more strongly suggestive of a causal effect of unilateral divorce laws than the TWFE estimate.
{adjustwidth}
-1.25in-1.25in A. Estimates of the effects of unilateral divorce laws TWFE ATT Callaway and Sant’Anna Wooldridge –0.604 –10.220 –5.530 (2.622) (3.086) (3.650) B. Internal validity of the TWFE estimand based on Proposition 5.3 uniformly in $\tau_{0}$ given $\tau_{0}$$\widehat{\mathbb{P}}(W^{*}=1)$ 0 0.3873 ^$\widehat{\mathbb{P}}(W^{*}=1\mid D=1)$ 0 0.6216 C. Internal validity of the TWFE estimand based on Proposition 5.4 uniformly in $\tau_{0}$ given $\tau_{0}$$\widehat{\mathbb{P}}(W^{*}=1)$ 0.1400 0.4802 ^$\widehat{\mathbb{P}}(W^{*}=1\mid D=1)$ 0.2246 0.7707
Table 1: Internal Validity of TWFE Estimand for Effects of Unilateral Divorce Laws
  • Notes: The data are a panel of the 1964–1996 U.S. The outcome is the state- and year-specific female suicide rate (per million women). The treatment is whether the state allowed unilateral divorce in a given year. The sample includes D.C. but excludes the states excluded by Goodman-Bacon (2021) as well as eight additional always-treated states. The measures of internal validity “given $\tau_{0}$” require an estimate of the CATE function, which we obtain using the approach of Wooldridge (2021).
While the TWFE estimate and the two estimates of the ATT are quite different, this paper focuses on another implication of the nonuniformity of the TWFE weight function. We ask: How representative of the underlying population is the TWFE estimand? What is the internal validity of this estimand if we are interested in the treated subpopulation? Panel B of Table 1 reports our estimates of $\mathbb{P}(W^{*}=1)$ and $\mathbb{P}(W^{*}=1\mid D=1)$, based on the representation of the TWFE estimand in de Chaisemartin and D’Haultfœuille (2020), revisited in our Proposition 5.3. First, because the weights on some group-time average treatment effects are negative, the TWFE estimand does not have a causal interpretation uniformly in $\tau_{0}$, ^$\widehat{\mathbb{P}}(W^{*}=1)=\widehat{\mathbb{P}}(W^{*}=1\mid D=1)=0$. Second, when we estimate the CATE function and use these estimates in constructing the bounds, we conclude that the TWFE estimand corresponds to the average treatment effect for at most 62.16% of the treated units or 38.73% of the entire population.
Panel C of Table 1 revisits these questions on the basis of the representation of the TWFE estimand in Proposition 5.4. Here, we assume that group-time average treatment effects are constant over time, which eliminates the problem of negative weights. Indeed, we now conclude that the TWFE estimand has a causal interpretation uniformly in $\tau_{0}$, even if it is still not particularly representative of the underlying population or the treated subpopulation. Our estimates of $\mathbb{P}(W^{*}=1)$ and $\mathbb{P}(W^{*}=1\mid D=1)$ are equal to 14.00% and 22.46%, respectively. When we use the estimated CATE function in constructing the bounds, these estimates increase to 48.02% and 77.07%. This is obviously much more than our initial estimate of 0, but still substantially less than 1, guaranteed in the case of $\mathbb{P}(W^{*}=1\mid D=1)$ when using the estimation methods in Callaway and Sant’Anna (2021), Wooldridge (2021), and other recent papers, each of which explicitly targets the ATT.
8   Conclusion
In this paper, we studied the representativeness and internal validity of a class of weighted estimands, which includes the popular OLS, 2SLS, and TWFE estimands in additive linear models. We examined the conditions under which such estimands can be written as the average treatment effect over a subpopulation. When a given estimand can be shown to correspond to the average treatment effect for a large subset of the population of interest, we say its internal validity is high. In our main results, we derived the sharp upper bound on the size of that subpopulation under different assumptions on treatment effect heterogeneity, which offers a practical tool to quantify the internal validity of weighted estimands.
Appendix A Proofs of Key Results
This appendix contains the proofs for our key results in Sections 3 and 4. Proofs for other results, including Theorem 4.2, can be found in the Supplemental Appendix.
Proof of Proposition 3.1.
To show the first claim, note that the equation $\mathbb{E}[W^{*}(Y(1)-Y(0))\mid W_{0}=1,X]=\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1,X]\cdot\underline{w}^{*}(X)$ holds since $W^{*}\perp\!\!\!\perp(Y(1),Y(0))\mid X,W_{0}=1$ by Definition 3.1. Since $\underline{w}^{*}(X)>0$, we can obtain
$\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1,X]$ $=\mathbb{E}[(Y(1)-Y(0))W^{*}\mid W_{0}=1,X]/\underline{w}^{*}(X)$ (A.1)
$=\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1,W_{0}=1,X]$
$=\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1,X],$
where the third equality holds from $W^{*}$ being a subpopulation of $W_{0}$.
The proposition’s second claim is established below:
$=\frac{\mathbb{E}[\underline{w}^{*}(X)\cdot\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1,X]\mid W_{0}=1]}{\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]}$
$=\mu(\underline{w}^{*},\tau_{0}).$
The first equality follows from $W^{*}$ being a subpopulation of $W_{0}$, the second from the law of iterated expectations and $W^{*}\leq W_{0}$, and the third from equation (A.1). ∎
Proof of Theorem 3.1.
()($\Longrightarrow$)( ⟹ ) First, suppose there exists $W^{*}\in\mathcal{W}(a;W_{0},\mathcal{T}_{\text{all}})$. By Proposition 3.1, we have that $\mu(a,\tau_{0})-\mu(\underline{w}^{*},\tau_{0})=0$ for all $\tau_{0}\in\mathcal{T}_{\text{all}}$. Let $\tau^{*}(X)=\frac{a(X)}{\mathbb{E}[a(X)\mid W_{0}=1]}-\frac{\underline{w}^{*}(X)}{\mathbb{P}(W^{*}=1\mid W_{0}=1)}$. We have $\mathbb{E}[a(X)^{2}]<\infty$ by Assumption 3.1 and $\mathbb{E}[\underline{w}^{*}(X)^{2}]\leq 1$, which holds by construction. Hence, $\mathbb{E}[\tau^{*}(X)^{2}]<\infty$, and thus $\tau^{*}\in\mathcal{T}_{\text{all}}$.
Thus, we must have $\mu(a,\tau^{*})-\mu(\underline{w}^{*},\tau^{*})=0$. Expanding this equality yields $0=\mu(a,\tau^{*})-\mu(\underline{w}^{*},\tau^{*})=\mathbb{E}[\tau^{*}(X)^{2}\mid W_{0}=1]$, which implies that
11 11 $=\mathbb{P}(\tau^{*}(X)=0\mid W_{0}=1)=\mathbb{P}\left(a(X)=\frac{\underline{w}^{*}(X)\mathbb{E}[a(X)\mid W_{0}=1]}{\mathbb{P}(W^{*}=1\mid W_{0}=1)}\mid W_{0}=1\right).$ (A.2)
Note that $\underline{w}^{*}(X)\geq 0$ and $\mathbb{P}(W^{*}=1\mid W_{0}=1)>0$ by construction. Also, $\mathbb{E}[a(X)\mid W_{0}=1]>0$ by Assumption 3.1. Therefore, $\mathbb{P}(a(X)\geq 0\mid W_{0}=1)=1$.